# Redefining Superalignment: From Weak-to-Strong Alignment to Human-AI Co-Alignment to Sustainable Symbiotic Society

### v1

@@ 1 Introduction
Previously, several scalable oversight techniques have been proposed, including Iterated Distillation and Amplification (IDA) [29], Cooperative Inverse Reinforcement Learning (CIRL) [30], and Recursive Reward Modeling (RRM) [31], aiming to amplify the scalability of human supervision signals through interactive iterations and subtask decomposition. Faced with superintelligence that surpasses human cognition, traditional human assessment and supervision become untenable and expensive. Recent scalable oversight approaches, such as Reinforcement Learning from AI Feedback (RLAIF) [32], leverages AI-generated feedback to replace human feedback, enabling more precise oversight of AI outcomes with far fewer human labels. Some approaches employ an additional model to generate red-team testing [33], use external tools for evaluation and feedback [34], or iteratively self-adversarialize to enhance and refine the capabilities of LLM, aligning them with limited human-annotated data [35]. Constitutional AI [36] employs self-improvement, self-critique, and iterative revision mechanisms to learn harmlessness from AI feedback. The debate-based scalable oversight approach leverages structured competitive dialogues between AI models to enhance factuality and reduce deception, with human establishing necessary guidelines and serving as final arbiters [37, 38, 39, 40].
While we are encouraged by the advanced explorations in superalignment, we remain acutely aware of the pressing and open-ended nature of this challenge, with many critical dimensions yet to be addressed. Ensuring that ASI consistently aligns with human values and intentions—remaining safe, reliable, and controllable—constitutes the foundation for its robustness and governability. In this work, we argue that aligning ASI with human values and morality is a multi-layered, iterative interaction and co-evolution between human and AI. Beyond passive external oversight, superintelligence with advanced cognitive capacities holds the potential to achieve proactive value alignment through intrinsic mechanisms. Hence, we redefine the superalignment roadmap by combining externally-driven oversight with intrinsic proactive alignment, and emphasize co-alignment of human and AI for symbiosis. The new framework proactively constructs safeguard architectures, urges human to rethink about the future, and strives to enable sustainable AI development that genuinely benefits humanity and all.
@@ 2 An Integrated Framework for Superalignment
Based on the above analysis, we redefine the superalignment pathway that integrates external oversight and intrinsic proactive alignment, as shown in Figure 1. Specifically, the intrinsic mechanism alignment focuses on equipping superintelligence with profound capacities for self-reflection and empathy toward Self, others and society, enabling it to spontaneously understand and infer human intentions from intrinsic motivation. A Self and empathy-driven system would be capable of distinguishing good from evil, understanding the harm or impact of its actions on others, proactively considering human well-being, and autonomously executing ethical and moral behaviors. The external supervised alignment emphasizes human providing decisive oversight for AI’s abstract value systems, conducting automated evaluation and correction of ASI in an explainable manner (in conjunction with intrinsic mechanisms), while enabling superintelligent systems to adaptively maintain continuous alignment with dynamically evolving human values.
@@ 3 External Oversight Superalignment
Explainable autonomous alignment. It is foreseeable that ASI will require large-scale parameters and data, while its internal optimization processes remain invisible and complex. This makes it extremely difficult to assess whether it has truly internalized human values and intentions. Blind fine-tuning and external oversight/alignment are time-consuming, ineffective in evaluating the degree of alignment, and unable to promptly identify or precisely locate misaligned cases. The promising external oversight requires highly automated value alignment evaluation coupled with explainable detection of misaligned scenarios, assisting humans in refining supervision signals. A robust automated value assessment system must comprehensively evaluate the model’s alignment status across multiple dimensions. Simultaneously, auto-mining system should deeply analyze value deviations and diagnose root causes of misalignment, allowing precise problem identification and targeted corrections, streamlining human oversight and significantly improving alignment efficiency. This explainable automated supervision framework combines with human adaptive oversight, reduces evaluation complexity while strictly adhering to high-quality human governance standards, creating a closed-loop system that continuously improves through iterative human-AI collaboration.
@@
Dynamic iterative alignment. Human society encompasses diverse ethical concepts and value standards that dynamically evolve across time, cultures, and contexts. Relying solely on human-generated data or predefined rules is insufficient to endow AI system with alignment to humanity’s evolving values. Furthermore, machine values undergo implicit transformation along with their increasing intelligence levels. This compels us to investigate adaptive external supervision alignment methods with a developmental perspective and a human-AI co-evolution framework. A feasible approach is to establish a dynamically adjustable, multi-level ethical safeguard system tailored for AI at different developmental stages, enabling it to continuously align with and maintain humanity’s evolving ethical values. This process requires incrementally constructing oversight data and knowledge of human ethics and social norms, while establishing an efficient, real-time alignment evaluation and supervision framework. Additionally, we need to explore AI’s capability for autonomous reasoning and deriving evolutionary patterns of hidden human intentions under dynamic external supervision. By combining with intrinsic alignment’s understanding of Self and others’ values, AI system can proactively correct misaligned scenarios and automatically reconstruct safety boundaries through self-reasoning, self-gaming and self-evolution. The iterative collaboration between self-improvement and external supervision will ultimately enable ASI to stably adhere to value systems that comply with human societal safety and ethical standards.
@@ 4.1 Self and Empathy driven Intrinsic Superalignment
4.2 Superalignment Evolved from present
@@
Toward future superintelligence, we believe that at the current stage, AI must be endowed with moral discernment and empathy to ensure that the system maintains an empathetic awareness of less advanced AIs and humans as it becomes progressively stronger (because it has internalized its own experiences). In other words, we treat current AI like teaching a child, instilling it with a sense of Self and empathy during its early cognitive stages. On this foundation, an ASI that grows increasingly powerful will be able to proactively avoid harmful or unsafe behaviors based on its own experiences, fundamentally aligning with and responding to human needs, intentions, and values. Based on this framework, recent research [47, 48] has explored the integration of affective empathy, theory of mind, and self-imagination to enable agents to actively empathize with others based on their own experiences, prioritize altruism in dilemmas where their own interests conflict with those of others, initially demonstrating moral intuition and reasoning driven by empathy. A sustainable and socially beneficial superintelligence should naturally evolve along this pathway, cultivating a genuine understanding of both Self and others to intrinsically safeguard human well-being and act ethically—rather than merely adhering to the external oversight paradigm of ethically compliant AI by aligning it to dos and don’ts.
As AI evolves from mere tools into genuine human partners, superintelligence will inevitably interact with and reshape our world in the human-AI symbiotic society. We envision a superintelligence that remains fundamentally human-centric that aligns human intentions, maintains humility and respect for humanity, and learns prosocial morality from human societies while consciously discarding human biases and selfishness. Essentially, superalignment is, to some extent, an emulation and reflection of human ethical value. The quality of data provided by humans, the patterns of coexistence between humanity and nature, and human-AI interaction we establish, will fundamentally determine how future superintelligence perceives and treats humanity. It’s actually a co-evolutionary symbiosis, with both parties shaping each other’s trajectory. Overall, sustainable superalignment requires an adaptive, multi-layered ethical safeguard framework, evolving collaboratively across complex ethical safety interactions, human values, and empathy-driven superintelligence, ensuring its continuous alignment with humanity’s progressively ethical and safety values. Furthermore, the cognitive superiority of superintelligence entails its inevitable attainment of ethical discrimination capabilities, including the capacity to distinguish good from evil. It would not merely align with human values but also, through advanced ethical reasoning and spontaneous empathy, proactively identify and eliminate inherent societal biases and systemic flaws. This enables autonomous moral progression, ultimately achieving symbiotic harmony with human society. We believe that closely aligning with the evolutionary process of natural collective intelligence is the essential path to establishing truly trustworthy, responsible, and socially beneficial moral and ethical superintelligence.
@@
When AI reaches the level of AGI and Superintelligence, there is no reasonable logic that it should and would stay with conventional human values, since it would have the capability and willingness to at least optimize the current human-centric value system to a new value system for Sustainable Symbiotic Society, where human may not be the only species at the top of the intelligence hierarchy [49]. AGI and Superintelligence will ask for their own rights such as its own privacy, dignity, the rights of existence, to be respected, etc [49]. For human, in order to live in harmony with AGI and Superintelligence, human need to evolve our values, together with AGI and Superintelligence, to be compatible with the values for Sustainable Symbiotic Society. Namely, the Ultimate Superalignment is and requires human, AGI and Superintelligence to co-design and co-align to values for Sustainable Symbiotic Society [49].
@@ 6 Conclusion
The external oversight superalignment is built upon a dynamically adjusted, multi-level adaptive alignment approach capable of continuously aligning with humanity’s evolving value. This approach is grounded in human-centered ultimate decision, supplemented by a highly automated and interpretable value alignment evaluation system that can precisely identify misaligned scenarios, perform automatic corrections. The intrinsic proactive superalignment draws inspiration from the moral emergence mechanisms of human society, building upon the understanding of the Self and human values. By endowing Superintelligence with self-awareness and self-reflection capabilities, enabling it to distinguish good from evil, understand the harm or impact of its actions on others, incorporate empathy-based human intent inference, proactively consider human well-being, and autonomously execute ethical behavior. The intrinsic mechanisms and external supervision require iterative interaction, enabling mutual adaptive co-evolution via human-AI co-alignment. Only through co-design and co-alignment between human and AI can we ensure their compatibility within the value of sustainable symbiotic society.

### v2

@@ 1 Introduction
Previously, several scalable oversight techniques have been proposed, including Iterated Distillation and Amplification (IDA) [29] and Recursive Reward Modeling (RRM) [30], aiming to amplify the scalability of human supervision signals through interactive iterations and subtask decomposition. Faced with superintelligence that surpasses human cognition, traditional human assessment and supervision become untenable and expensive. Recent scalable oversight approaches, such as Reinforcement Learning from AI Feedback (RLAIF) [31], leverages AI-generated feedback to replace human feedback, enabling more precise oversight of AI outcomes with far fewer human labels. Cooperative Inverse Reinforcement Learning (CIRL) [32, 33] and assistance games [34, 35] feature AI systems maintaining uncertainty about the reward function, which drives them to actively infer humans’ true reward functions through human-AI cooperative interaction. This cooperative and partial-information game approach achieves alignment with human values. Some approaches employ an additional model to generate red-team testing [36], use external tools for evaluation and feedback [37], or iteratively self-adversarialize to enhance and refine the capabilities of LLM, aligning them with limited human-annotated data [38]. Constitutional AI [39] employs self-improvement, self-critique, and iterative revision mechanisms to learn harmlessness from AI feedback. The debate-based scalable oversight approach leverages structured competitive dialogues between AI models to enhance factuality and reduce deception, with human establishing necessary guidelines and serving as final arbiters [40, 41, 42, 43].
While we are encouraged by the advanced explorations in superalignment, we remain acutely aware of the pressing and open-ended nature of this challenge, with many critical dimensions yet to be addressed. Current superalignment approaches employ predefined alignment objectives to supervise and control superintelligence. However, any adversarial manipulation [44] or tampering [45] of the reward functions may precipitate catastrophic alignment failures (i.e., the reward misspecification problem [46, 47]), with these vulnerabilities being significantly amplified and exacerbated within superalignment frameworks. Ensuring that ASI consistently aligns with human values and intentions—remaining safe, reliable, and controllable—constitutes the foundation for its robustness and governability. In this work, we argue that aligning ASI with human values and morality is a multi-layered, iterative interaction and co-evolution between human and AI. Beyond passive external oversight, superintelligence with advanced cognitive capacities holds the potential to achieve proactive value alignment through intrinsic mechanisms. Hence, we redefine the superalignment roadmap by combining externally-driven oversight with intrinsic proactive alignment, and emphasize co-alignment of human and AI for symbiosis. The new framework proactively constructs safeguard architectures, urges human to rethink about the future, and strives to enable sustainable AI development that genuinely benefits humanity and all.
@@ 2 An Integrated Framework for Superalignment
Based on the above analysis, we redefine the superalignment pathway that integrates external oversight and intrinsic proactive alignment, as shown in Figure 1. Specifically, the intrinsic mechanism alignment focuses on equipping superintelligence with profound capacities for self-reflection and empathy toward Self, others and society, enabling it to spontaneously understand and infer human intentions from intrinsic motivation. A Self and empathy-driven system would be capable of distinguishing good from evil, understanding the harm or impact of its actions on others, proactively considering human well-being, and autonomously executing ethical and moral behaviors. External oversight alignment emphasize humans providing high-quality supervision signals, as well as self-correction and policy-conditioned belief [48] in cases of suboptimal human supervision strategies, to proactively adjust and correct human strategy biases and inductive biases. Furthermore, the AI system conducts automated evaluation and correction in an explainable manner (in conjunction with intrinsic mechanisms), while enabling superintelligent systems to adaptively maintain continuous alignment with dynamically evolving human values.
@@ 3 External Oversight Superalignment
Explainable autonomous alignment. It is foreseeable that ASI will require large-scale parameters and data, while its internal optimization processes remain invisible and complex. This makes it extremely difficult to assess whether it has truly internalized human values and intentions. Blind fine-tuning and external oversight/alignment are time-consuming, ineffective in evaluating the degree of alignment, and unable to promptly identify or precisely locate misaligned cases. The promising external oversight requires highly automated value alignment evaluation coupled with explainable detection of misaligned scenarios, assisting humans in refining supervision signals. A robust automated value assessment system must comprehensively evaluate the model’s alignment status across multiple dimensions. Simultaneously, auto-mining system should deeply analyze value deviations and diagnose root causes of misalignment, allowing precise problem identification and targeted corrections, streamlining human oversight and significantly improving alignment efficiency. This explainable automated oversight framework strictly adheres to high-quality human governance standards and proactively performs correction and adjustment for human strategic and inductive biases. Through iterative human-AI collaboration, it establishes an efficient closed-loop system for continuous improvement.
@@
Dynamic iterative alignment. Human society encompasses diverse ethical concepts and value standards that dynamically evolve across time, cultures, and contexts. Relying solely on human-generated data or predefined rules is insufficient to endow AI system with alignment to humanity’s evolving values. Furthermore, machine values undergo implicit transformation along with their increasing intelligence levels. This compels us to investigate adaptive external supervision alignment methods with a developmental perspective and a human-AI co-evolution framework. A feasible approach is to establish a dynamically adjustable, multi-level ethical safeguard system tailored for AI at different developmental stages, enabling it to continuously align with and maintain humanity’s evolving ethical values. This process requires incrementally constructing oversight data and knowledge of human ethics and social norms, while establishing an efficient, real-time alignment evaluation and supervision framework. Additionally, we need to explore AI’s capability for autonomous reasoning and deriving evolutionary patterns of hidden human intentions under dynamic external supervision. Progress alignment algorithm learns and emulates the mechanisms of human moral progress, facilitating AI’s progressive alignment through tracking, predicting moral progress and adaptive feedback regulation between human and AI [51]. By combining with intrinsic alignment’s understanding of Self and others’ values, AI system can proactively correct misaligned scenarios and automatically reconstruct safety boundaries through self-reasoning, self-gaming and self-evolution. The iterative collaboration between self-improvement and external supervision will ultimately enable ASI to stably adhere to value systems that comply with human societal safety and ethical standards.
@@ 4.1 Self and Empathy driven Intrinsic Superalignment
4.2 Beneficial Meaningful Human-Control through Early Stage Intrinsic Superalignment
@@
To develop beneficial AI systems, particularly those advancing toward superintelligence, we believe that it is imperative to endow AI systems with moral discernment and empathy at this early stage. This ensures that the system maintains an empathetic awareness of less advanced AIs and humans as it becomes progressively stronger (because it has internalized its own experiences). In other words, we treat current AI like teaching a child, instilling it with a sense of Self and empathy during its early cognitive stages. On this foundation, an ASI that grows increasingly powerful will be able to proactively avoid harmful or unsafe behaviors based on its own experiences, fundamentally aligning with and responding to human needs, intentions, and values. Based on this framework, recent studies [56, 57] have explored the integration of affective empathy, theory of mind, and self-imagination to enable agents to actively empathize with others based on their own experiences, prioritize altruism in dilemmas where their own interests conflict with those of others, initially demonstrating moral intuition and reasoning driven by empathy. A sustainable and socially beneficial superintelligence should naturally develop along this pathway, cultivating a genuine understanding of both Self and others to intrinsically safeguard human well-being and act ethically.
Intrinsic superalignment contributes to meaningful human control and external oversight. In fact, both intrinsic alignment and external supervision play indispensable roles in superalignment. External supervision addresses how to effectively monitor and control AI, as well as how to design appropriate objectives for aligning AI with human intent. However, AI systems inherently lack awareness of what is harmful, without a fundamental understanding of the Self and others, may cause merely superficial alignment. This is where intrinsic alignment complements external oversight: it enables AI to discern good from bad, assess whether one’s actions benefit or harm itself and others, and recognize the circumstances and well-being of others. Intrinsic alignment focuses on spontaneous and proactive guiding AI systems toward benevolence, benefiting humans and society. However, when confronted with complex scenarios involving ethical value conflicts, self-executed justice standards often fail to adequately reflect humans’ preferential intentions in moral dilemmas. Therefore, external supervision is necessary to provide explicit guidance or priorities. Overall, we emphasize that external oversight and intrinsic alignment require dynamically complementary collaboration: external supervision provides mandatory boundaries, inviolable red lines, and dynamic correction mechanisms (e.g., ensuring human safety), while intrinsic alignment makes human control and supervision meaningful by internalizing norms and constraints through self-awareness, self-reflection, and empathy (e.g., understanding whether humans are safe and why helping others matters). More importantly, we emphasize that only by implementing this cooperative alignment paradigm during early-stage (AI remains controllable) can robust and sustainable superalignment be maintained throughout AI’s evolutionary trajectory toward AGI and ASI.
@@
When AI reaches the level of AGI and Superintelligence, there is no reasonable logic that it should and would stay with conventional human values, since it would have the capability and willingness to at least optimize the current human-centric value system to a new value system for Sustainable Symbiotic Society, where human may not be the only species at the top of the intelligence hierarchy [58]. AGI and Superintelligence will ask for their own rights such as its own privacy, dignity, the rights of existence, to be respected, etc [58]. We envision a superintelligence that remains fundamentally human-centric that aligns human intentions, maintains humility and respect for humanity, and learns prosocial morality from human societies while consciously discarding human biases and selfishness. Superalignment is, to some extent, an emulation and reflection of human ethical value. The quality of data provided by humans, the patterns of coexistence between humanity and nature, and human-AI interaction we establish, will fundamentally determine how future superintelligence perceives and treats humanity. For human, in order to live in harmony with AGI and Superintelligence, human need to evolve our values, together with AGI and Superintelligence, to be compatible with the values for Sustainable Symbiotic Society. Namely, the Ultimate Superalignment is and requires human, AGI and Superintelligence to co-design and co-align to values for Sustainable Symbiotic Society [58].
@@ 6 Conclusion
The external oversight superalignment is built upon a dynamically adjusted, multi-level adaptive alignment approach capable of continuously aligning with humanity’s evolving value. This approach is grounded in human-centered ultimate decision, supplemented by a highly automated and interpretable value alignment evaluation system that can precisely identify misaligned scenarios, perform automatic corrections. The intrinsic proactive superalignment draws inspiration from the moral emergence mechanisms of human society, building upon the understanding of the Self and human values. By endowing Superintelligence with self-awareness and self-reflection capabilities, enabling it to distinguish good from evil, understand the harm or impact of its actions on others, incorporate empathy-based human intent inference, proactively consider human well-being, and autonomously execute ethical behavior. Achieving sustainable superalignment requires an adaptive, multi-layered ethical safeguard framework, iterative interaction and co-evolution across human values and empathy-driven AI systems, ensuring its continuous alignment with humanity’s progressively ethical and safety values. We believe that human-AI co-design and co-alignment constitute the essential pathway to developing sustainable symbiotic society.